Consider Snackistan, a hypothetical country that produces only cakes. In 2016, a cake is priced at $1.00. Complete the first row of the table with the quantity of cakes that can be bought with $100. Hint: In this problem, assume it is not possible to buy a fraction of a cake, and always round down to the nearest whole cake. For example, if your calculations result in 1.5 cakes, the answer should be 1 cake. Year 2016 2017 Price of a Cake (Dollars) 1.00 Cakes Bought with $100 (Quantity) Suppose the government of Snackistan cannot raise sufficient tax revenue to pay its debts. In order to meet its debt obligations, the government prints money. As a result, the money supply rises by 30% by 2017. Assuming monetary neutrality holds, complete the second row of the table with the new price of a cake and the new quantity of cakes that can be bought with $100 in 2017. The impact of the governments decision to raise revenue by printing money on the value of money is known as the Grade It Now Save & Continue Continue without saving
Answer #1

Year

Price of a cake

(in $)

Cakes bought with $100

(Quantity)

2016

1.00

100

2017

1.20

76

2016

Quantity that can be purchased: $100/$1 = 100

2017

If money supply rises by 30% and neutrality of money holds then price level rises by 30%

New price = 1 + (1x30%) = 1 + (1x30/100) = 1+0.30 = $1.30

Quantity that can be purchased: $100/$1.30 = 76.92 (Round down) = 76

The impact of the government's decision to raise revenue by printing money on the value money is known as Inflation Tax

  • Inflation tax -The value of the money will decline.
  • The revenue raised by inflation tax is called seigniorage